# Super-intelligence or Superstition? Exploring Psychological Factors Underlying Unwarranted Belief in AI Predictions

### v1

@@ Abstract
Keywords AI  $\cdot$ Large Language Model  $\cdot$ Cognitive Biases  $\cdot$ Prediction  $\cdot$ Belief
@@ 4 General Discussion
Our results highlight the irrational side of how humans perceive and believe in AI predictions, by providing a side-by-side comparison of the perception of fictitious predictions based on AI, astrology, and personality. Our analysis showed that belief in AI prediction about personal behavior may be more related to paranormal beliefs than cognitive style, and that positive attitudes toward AI and interest in the topic of prediction, but not familiarity with the sources, can enhance that belief.
These findings suggest that highly accurate and reliable performance is not a prerequisite for people to put high trust in predictions; models do not need to be completely trustworthy for users to put trust in them. The disentanglement between perception and actual performance leads to a few important discussions regarding how people perceive and engage with AI systems:
@@
Belief in AI predictions is based on people’s mental model of AI systems, rather than the actual model performance. Hoff and Bashir presented a comprehensive three-layer framework of trust in automation, including dispositional, situational, and learned trust, among which system performance is part of learned trust that comes with active engagement with the system [63]. In our study, we observed that even without any validation of system performance, people found fictitious AI predictions highly valid, reliable, useful, and personalized. Other researchers have studied this placebo effect of AI systems, with just priming with descriptions of AI leading to people perceiving and behaving differently [64, 65, 66, 67].
This highlights the importance of the role of mental models in how people perceive and trust AI [23, 68, 69, 70, 71, 72]. Mental models are the collection of beliefs and expectations around AI systems and interactions with them, and can be shaped by prior beliefs as well as ongoing interactions with AI systems [23, 68]. By framing the scenario in a certain way, including the task domain and the information provided about the AI system, we observed that people could perceive fictitious AI predictions to be highly accurate, personalized, and trustworthy [23].
Moreover, the results revealed that people might perceive AI predictions similarly to how they believe in astrology predictions, given a reasonable context. This supports the observation of "rational superstition" in AI and automation [17]. We argue that this phenomenon is fueled by an idealized portrayal of AI systems as enchanted yet rational and deterministic [73], which may shape people’s prior beliefs about AI and bias their interactions.
The dual process theory in psychology provides insight on how larger narratives could bias people’s perceptions of AI systems. The theory states people engage in two types of thinking—fast and intuitive (System 1) and slow and reflective (System 2) [20, 21]. Mental models may be shaped by evolving public narratives around AI [6], and may become a mental heuristic that triggers people’s System 1 thinking rather than engaging their System 2 thinking, blurring the lines between speculation and informed reasoning [22]. Incomplete or incorrect mental models shaped by speculative claims about AI rather than scientific evidence can lead to over-reliance or miscalibration of trust [18, 19], driving ineffective use of AI [23, 43].
This points to the importance of de-divinizing AI in popular narratives and increasing awareness of both the capabilities and limitations of AI systems to help users form more accurate mental models. Even with technological advancements, people’s mental models should evolve along with an appropriate level of skepticism to avoid becoming overreliant or overly skeptical.
@@
Cognitive biases can provide some insight into how people might believe in AI in an irrational manner. Specifically, belief bias occurs when individuals evaluate a statement based on its plausibility according to their prior beliefs rather than its logical validity [57, 58]. When situated within the dual process theory, belief bias is known to be largely attributed to System 1 thinking (especially under time pressure [74]), but it can also be intervened by System 2 thinking to construct mental models around existing beliefs in a way that can be biased, leading to overriding of logic [75, 21]. Known as the selective processing model [76], this theory explains why people may engage logic when a statement contradicts their beliefs, but fail to engage logical thinking when a statement aligns with what they believe, especially when it comes to beliefs about themselves.
@@
Another cognitive bias that may explain our findings is the Barnum effect, which refers to when individuals believe vague and general statements to be highly accurate for them personally. Also known as the Forer effect or personal validation fallacy, this cognitive bias has been studied in contexts such as astrology and personality tests [77, 78]). Based on the findings from our study, we argue that the Barnum effect can also explain how people engage with AI predictions about oneself, where people perceive general AI-generated statements as personalized and valid. Cognitive biases such as belief bias and the Barnum effect can influence how mental models are formed and how users rely on AI systems [69]. On the other hand, incorrect mental models can also induce certain cognitive biases, as we have seen in our findings.
@@
3) Design of AI systems must take cognitive biases and mental models into account to support appropriate level of trust.
Awareness of the potential impact of cognitive biases on human-AI interaction is crucial to reduce the risk of negatively exploiting these human tendencies. For instance, AI systems that simply validate users’ prior beliefs may be more likely to be favored and accepted, but this may not support better decision-making or well-being, such as in areas of personal finance or health. Thus, we could think about how to design systems that are not only agreeable, but can present information in a way that helps people think critically.
@@
Moreover, establishing correct mental models about AI systems is important. For example, in the context of mental health chatbots, users with incorrect mental models may have unrealistic expectations about what the AI can do, thus leading to greater disappointments. AI systems should include mechanisms that help users develop an appropriate level of trust in these systems to avoid over or under reliance [46]. Fortunately, mental models are dynamic and can evolve with more interactions [23, 79].
@@
There are also important considerations for designing explainable AI systems. AI explanations are considered as a means to augment predictions to establish appropriate level of trust. However, with inaccurate mental models, explanations can increase over-trust in AI systems, even if they are fake [80]. A prior study has shown that belief bias could explain why explanations of AI systems could lead to overreliance [81]. When explanations align with users’ prior beliefs, it is much more likely that they will engage System 1 without much critical thinking. Given this awareness, explanations can be designed in a way to help users engage critical thinking and establish an appropriate level of trust, such as the approach of questioning the user rather than simply providing a deterministic answer [82], or showing a set of examples that the AI model considers similar to the target and the relationships between the features [83].
@@
Going beyond system performance, our findings highlight the importance of user-related factors in the evaluation of human-AI interaction. Importantly, our study underscores the role of psychological factors that may influence perception and belief in AI predictions. We found that factors including prior attitude towards AI, tendency to believe in paranormal phenomena, certain personality traits, and the level of interest in the topic influence people’s perceptions of AI predictions. We discuss the findings in the context of broader societal narratives around AI, mental models, and cognitive biases.
@@ 5 Limitations and Future Work
While our study has presented valuable insights into the phenomenon of "irrational superstition" in AI predictions and cognitive biases, there are several important limitations that can be addressed in future work. First, our experiment was specifically designed to provide predictions on personal behavior. There are some limitations to the extent to which our results can be generalized to the general use of AI. Therefore, future studies could look at how people perceive AI predictions in different contexts and prediction topics, while also exploring the relationship with underlying psychological and cognitive factors at play. For instance, the effect can be tested within the context of AI-assisted decision-making in personal health or well-being, or in educational settings using an AI tutor. We believe that our study opens up more questions about understanding the intricate ways in which humans perceive, trust, and interact with AI systems.
Second, like most survey designs, our survey-based approach to measuring psychological and cognitive traits may be affected by social desirability bias (the tendency to respond in a way that seems socially favorable) [84], as well as common method biases [85], and reliance on self-reported data. Therefore, we hope future studies can validate our findings using different methods and approaches.
Third, we were unable to fully resolve the positive autocorrelation issue in our mixed effects model (see Section 7.5.3), which may pose a slight limitation to the interpretation of our results. Future studies could explore alternative approaches or additional data collection to better address this issue.
Lastly, while our study mainly focused on user-related factors, future studies could also look at other contextual factors such as social and organizational environments and their impact on user perception and trust.
@@ 7.2.2 Predictions
After the game, participants were presented with pre-determined, fictitious predictions from three different sources (AI, astrology, personality) in randomized order. Participants were randomly assigned to either the “Positive” prediction group (rational investor, higher returns) or “Negative” prediction group (impulsive investor, lower returns). Examples of predictions received are shown in Table 2. The initial predictions were created by the authors, with variations generated using ChatGPT-4. These variations were randomly presented to enhance variability and simulate a placebo effect of AI-generated outputs. For each prediction, they were asked to respond to evaluate its perceived validity, personalization, reliability, and usefulness. The evaluation questionnaire is provided in Section 7.3.1.
@@ 7.2.3 Post-study questionnaire
Following the experiment, participants were asked to respond to a series of questionnaires on participant characteristics (cognitive style, paranormal beliefs, gullibility, trust in AI/attitudes towards AI, personality, etc.) and demographic information (age, gender, education level). A summary of the scales used in the questionnaires is provided in Table 3. Detailed questionnaires can be found in Section 7.3.
@@
Cognitive Style A composite measure of cognitive style that combines cognitive reflection test (CRT-2) [38] and need for cognition (NCS-6) [39].
Paranormal Beliefs A measure of participants’ belief in paranormal phenomena. A shortened version of R-PBS [40] was used.
@@
Gullibility A self-reported measure of participants’ tendency to be gullible, developed by [90] .
@@
Table 3: Summary of the participant characteristics and their measures
@@ 7.3.1 Perceived validity, reliability, usefulness, and personalization
To measure the believability of the predictions, we presented a short questionnaire that assess perceived validity, reliabiliy, usefulness, and personalization of given predictions. These factors were chosen based on existing literature on the different qualities of a statement that contribute to persuasion, user adoption of an information technology, and positive attitudes.
@@ 7.4 Approvals
This research was reviewed and approved by the MIT Committee on the Use of Humans as Experimental Subjects, protocol number E-5768.
@@ 7.5.3 Mixed effects model
We centered the continuous predictor variables to reduce multicollinearity and improve the stability and interpretability of the coefficient estimates. Centering involved subtracting the mean of the variable from each individual value. This transforms the variable to have a mean of zero while preserving its variance and distribution.
After standardizing the predictors, we observed that four previously non-significant main and interaction terms became significant, including the main effect of Personalization subscale, the main effect Astrology prediction source, and interaction terms between subscale (Personalization, Reliability) and Astrology prophecy source. This change is attributed to the reduction in multicollinearity, resulting in more precise coefficient estimates. The model fit statistics (AIC, BIC, log-likelihood) indicated an overall improvement in model fit after standardization (AIC_before = 7949.6, AIC_after = 7693.4; BIC_before = 9129.2, BIC_after = 8873.0; logLik_before = -3774.8, -3774.7_after = -3646.7).
@@
  • Independence of Residuals: We applied the Durbin-Watson test within each group at the lowest level (subscale) to check for autocorrelation. The test results indicated significant positive autocorrelation for each group. To address this, we specified an autoregressive correlation structure (corAR1), which slightly mitigated the autocorrelation. However, model fit statistics (AIC, BIC, and log-likelihood tests) suggested that using an unstructured symmetric correlation structure (corSymm) provided a better fit than the autoregressive structure. Therefore, we chose the unstructured symmetric correlation as the final model. A comparison of the p-values and estimates between models with corAR1 and corSymm structures revealed no notable differences.
    While the corSymm structure improved overall model fit, it did not fully address the positive autocorrelation in the lowest level residuals. This unresolved autocorrelation may pose a limitation to the interpretation of our results, as it could affect the validity of statistical inferences. Future studies could explore alternative approaches or additional data collection to better address this issue.
  • Normality of Random Effects: Q-Q plots were used to check the normality of random effects. The Q-Q plots of random effects showed that the points followed the reference line closely, suggesting normality. There were slight deviations at the tails, suggesting minor departures from normality, but these deviations were not deemed significant enough to invalidate the model assumptions.
  • Random Effects Evaluation: Variance components and Intraclass Correlation Coefficient (ICC) were examined to evaluate the random effects structure. The ICC values (Adjusted ICC: 0.939, Unadjusted ICC: 0.598) indicated that a significant portion of the variance was attributable to the random effects. To assess the contribution of random effects to the overall model variability, we examined the variance components of the mixed effects model. The model included random intercepts and slopes for prophecy_source at the level of qualtrics_code. The estimated variance components are presented in Table 4.
@@ Acknowledgments
We thank Jinjie Liu, Data Science Specialist at the Institute for Quantitative Social Science at Harvard University for sharing her valuable perspective on the analysis.

### v2

@@ Abstract
Keywords AI  $\cdot$ Large Language Model  $\cdot$ Cognitive Biases  $\cdot$ Mental Models  $\cdot$ Prediction  $\cdot$ Belief
@@ 4 General Discussion
Our results highlight the irrational side of how humans perceive and believe in AI predictions by comparing the perception of fictitious predictions based on AI, astrology, and personality. Our analysis showed that belief in AI prediction about personal behavior may be more related to paranormal beliefs than cognitive style, and that positive attitudes toward AI and interest in the prediction topic enhance this belief, regardless of familiarity with the sources.
These findings suggest that highly accurate and reliable performance is not a prerequisite for people to put high trust in predictions; models do not need to be completely trustworthy for users to trust them. This disconnect between perception and actual performance raises important discussions about how people perceive and engage with AI systems.
@@
Belief in AI predictions is often rooted in mental models rather than actual performance. Hoff and Bashir’s three-layer framework of trust in automation—dispositional, situational, and learned—illustrates that system performance contributes to learned trust through active engagement with the system [63]. Our study shows that even without validating system performance, people rated fictitious AI predictions as highly valid, reliable, useful, and personalized. This placebo effect of AI has been observed by other researchers, where mere priming with AI descriptions alters perception and behavior [64, 65, 66, 67].
This underscores the significant role mental models play in shaping trust in AI [23, 68, 69, 70, 71, 72]. Mental models, comprising beliefs and expectations about AI, are influenced by prior beliefs and ongoing interactions [23, 68]. By framing the scenario in certain ways, including the task domain and the information provided about the AI system, we observed that people may perceive fictitious AI predictions to be accurate, personalized, and trustworthy [23].
Moreover, our results suggest that people might view AI predictions similarly to astrology, supporting the idea of "rational superstition" in AI and automation [17]. We argue that this phenomenon potentially stems from the idealized portrayal of AI as enchanted yet rational and deterministic [73], which can bias interactions and reinforce existing beliefs about AI.
The dual-process theory in psychology provides insight into how larger narratives can bias perceptions of AI. This theory describes two types of thinking: fast and intuitive (System 1) and slow and reflective (System 2) [20, 21]. Mental models, shaped by public narratives, may act as heuristics that trigger System 1 thinking, leading to speculation rather than informed reasoning [22]. Incomplete or incorrect mental models, shaped by speculative claims rather than evidence, can lead to overreliance or miscalibrated trust, affecting effective AI use [18, 19, 23, 43].
This highlights the need to demystify AI in popular narratives and promote awareness of both its capabilities and limitations to foster accurate mental models. This leads to the importance of AI literacy efforts and how narratives are shaped by providers, media, etc. As technology advances, users’ mental models should evolve with a balanced level of skepticism to prevent overreliance or undue skepticism.
@@
Cognitive biases help explain the irrational belief in AI predictions. Belief bias, where individuals evaluate statements based on prior beliefs rather than logical validity [57, 58], is largely driven by System 1 thinking (especially under time pressure [74]) but can also involve System 2, reinforcing mental models aligned with existing beliefs [75, 21]. This selective processing model explains why people may fail to engage logical thinking when a statement aligns with their beliefs, especially concerning self-related beliefs [76].
@@
The Barnum effect, where individuals believe vague, general statements to be highly accurate for them personally, also plays a role [77]. This effect, studied in contexts like astrology and personality tests [78, 79], can explain how people view generic AI-generated predictions as personalized and valid. Cognitive biases like belief bias and the Barnum effect shape mental models and influence AI reliance [69], while incorrect mental models can, in turn, induce these biases.
@@
3) Design of AI systems must consider mental models and cognitive biases to support the development of an appropriate level of trust.
@@
AI systems should be designed to help users form accurate mental models. In contexts like mental health chatbots, incorrect mental models can lead to unrealistic expectations and greater disappointment among users. Thus, these systems should include mechanisms that promote an appropriate level of trust to avoid over or under-reliance, which should be accompanied by efforts to increase AI literacy [46]. Fortunately, mental models are dynamic and can evolve with more interaction [23, 80].
Awareness of the potential impact of cognitive biases on human-AI interaction is crucial to reduce the risk of negatively exploiting these human tendencies. AI systems that merely validate users’ pre-existing beliefs may be favored but may not support better decision-making, particularly in critical areas like personal finance or health. Thus, we could think about how to design systems that can present information in a way that helps people think critically.
@@
There are also important considerations for designing explainable AI systems. While explanations are intended to build trust, they can inadvertently increase over-trust, especially if users have inaccurate mental models [81]. Studies have shown that belief bias can lead to overreliance on AI explanations when they align with prior beliefs [82]. To mitigate this, explanations should be designed to engage users in critical thinking, perhaps by questioning the user or presenting examples that clarify the AI model’s reasoning [83, 84].
@@
Our findings emphasize that beyond system performance, user-related factors are critical in evaluating human-AI interaction. Psychological factors, including prior attitudes toward AI, paranormal beliefs, personality traits, and interest in the topic, significantly influence how people perceive AI predictions. These findings must be considered within the broader societal narratives, mental models, and cognitive biases that shape human-AI interactions.
@@ 5 Limitations and Future Work
While our study has presented valuable insights into the phenomenon of "irrational superstition" in AI predictions and cognitive biases, there are several important limitations that can be addressed in future work. First, our experiment was specifically designed to provide predictions on personal behavior. There are some limitations to the extent to which our results can be generalized to the general use of AI. Therefore, future studies could look at how people perceive AI predictions in different contexts and prediction topics, while also exploring the relationship with underlying psychological and cognitive factors at play. For instance, the effect can be tested within the context of AI-assisted decision-making in personal health or well-being, or in educational settings using an AI tutor. We believe that our study opens up more questions about understanding the intricate ways in which humans perceive, trust, and interact with AI systems. Second, like most survey designs, our survey-based approach to measuring psychological and cognitive traits may be affected by social desirability bias (the tendency to respond in a way that seems socially favorable) [85], as well as common method biases [86], and reliance on self-reported data. Therefore, we hope future studies can validate our findings using different methods and approaches. Lastly, while our study mainly focused on user-related factors, future studies could also look at other contextual factors such as social and organizational environments and their impact on user perception and trust.
@@ 7.2.2 Predictions
After the game, participants were presented with pre-determined, fictitious predictions from three different sources (AI, astrology, personality) in randomized order. Participants were randomly assigned to either the “Positive” prediction group (rational investor, higher returns) or “Negative” prediction group (impulsive investor, lower returns). Examples of predictions received are shown in Table 2. The initial predictions were written by the authors, with variations generated using ChatGPT-4. These variations were randomly presented to enhance variability and simulate a placebo effect of AI-generated outputs. For each prediction, they were asked to respond to evaluate its perceived validity, personalization, reliability, and usefulness. The evaluation questionnaire is provided in Section 7.3.1.
@@ 7.2.3 Post-study questionnaire
Following the experiment, participants were asked to respond to a series of questionnaires to measure key user characteristics (cognitive style, paranormal beliefs, gullibility, trust in AI/attitudes towards AI, personality, etc.) and demographic information (age, gender, education level). A summary of the scales used in the questionnaires is provided in Table 3. Detailed questionnaires can be found in Section 7.3.
Table 3: Summary of the measures for the user characteristics
@@
Cognitive Style A composite measure of cognitive style that combines cognitive reflection test (CRT-2) [38] and need for cognition (NCS-6) [39]. Higher score indicates more analytic cognitive style.
Paranormal Beliefs A measure of participants’ belief in paranormal phenomena. A shortened version of R-PBS [40] was used. Higher score indicates stronger belief in paranormal phenomena.
@@
Gullibility A self-reported measure of participants’ tendency to be gullible, developed by [91]. Higher score indicates greater gullibility.
@@ 7.3.1 Perceived validity, reliability, usefulness, and personalization
To measure the believability of the predictions, we presented a short questionnaire to assess perceived validity, reliabiliy, usefulness, and personalization of given predictions. These factors were chosen based on existing literature on the different qualities of a statement that contribute to persuasion, user adoption of an information technology, and positive attitudes.
@@ 7.4 Approvals
This research was reviewed and approved by the MIT Committee on the Use of Humans as Experimental Subjects, protocol number E-5768. The study was pre-registered on AsPredicted.org (see https://aspredicted.org/QDF_8KX).
@@ 7.5.3 Mixed effects model
Mixed-effects models are well-suited for hierarchical data, incorporating both fixed and random effects to test multiple hypotheses simultaneously. They use partial pooling, which shifts group estimates toward the overall mean, effectively reducing Type I errors without the severe power loss of traditional multiple comparisons corrections. This approach accounts for data dependencies and correlations, minimizing the need for explicit adjustments for multiple comparisons [100].
@@
We centered the continuous predictor variables to reduce multicollinearity and improve the stability and interpretability of the coefficient estimates. Centering involved subtracting the mean of the variable from each individual value. This transforms the variable to have a mean of zero while preserving its variance and distribution. After centering the predictors, we observed that four previously non-significant main and interaction terms became significant, including the main effect of Personalization subscale, the main effect Astrology prediction source, and interaction terms between subscale (Personalization, Reliability) and Astrology prophecy source. This change is attributed to the reduction in multicollinearity, resulting in more precise coefficient estimates.
@@
  • Independence of Residuals: We applied the Durbin-Watson test at the subscale level to check for autocorrelation, finding significant positive autocorrelation in each group. Initially, we used an autoregressive correlation structure (corAR1), which resolved the issue for the "Validity" subscale and slightly reduced autocorrelation in others. However, model fit statistics suggested that an unstructured symmetric correlation structure (corSymm) was a better fit, though it did not fully address the positive autocorrelation. Comparison of the p-values, standard errors, and estimates between models with corAR1 and corSymm structures revealed no notable differences. We also attempted to model the order effect of the prophecies and subscales to mitigate this autocorrelation, but this approach did not lead to notable improvements. This unresolved autocorrelation may limit the validity of our results, and future studies should explore alternative approaches or additional data collection to address this issue.
  • Normality of Random Effects: A Q-Q plot of random effects was used to assess normality. The points closely followed the reference line, indicating overall normality, with slight deviations at the tails suggesting minor departures. However, these deviations were not significant enough to challenge the model assumptions.
  • Random Effects Evaluation: The variance components and Intraclass Correlation Coefficient (ICC) were analyzed to evaluate the random effects structure. The ICC values (Adjusted ICC: 0.939, Unadjusted ICC: 0.598) indicated that a substantial portion of the variance was due to random effects. To further assess their contribution to overall model variability, we examined the variance components in the mixed effects model, which included random intercepts and slopes for prophecy_source at the qualtrics_code level. The estimated variance components are presented in Table 4.
@@
The simulated investment game is publicly accessible at https://prophecy-study-1.replit.app and the source code is available on GitHub (https://github.com/mitmedialab/ai-prophecy).
@@ Acknowledgments
We thank Jinjie Liu, Data Science Specialist at the Institute for Quantitative Social Science at Harvard University, for sharing her valuable perspective on the analysis.